FOURIER TRANSFORM MICROWAVE SPECTRUM OF MgCCH (X^{2}Σ^{+})

J. MIN, D. T. HALFEN, M. SUN, B. T. HARISS, L. M. ZIURYS, University of Arizona, Deptment of Chemistry and Biochemitry and Steward Observatory, Tucson, AZ-85721; D. J. CLOUTHIER, University of Kentucky, Deptment of Chemistry, Lexington, KY-40506.

The pure rotational spectrum of MgCCH (X^{2}Σ^{+}) in the frequency range of 9-40 GHz has been measured using Fourier transform microwave (FTMW) methods. The molecule was synthesized using discharge assisted laser ablation spectroscopy (DALAS) from a mixture of 0.1% acetylene in argon and the ablation of a magnesium rod. From these data, the hydrogen hyperfine parameters have been determined for the first time, as well as refinement of the rotational and spin-rotational constants, combined with previous millimeter-wave spectra measured by the Ziurys group.